please help me solve this question
122 liters to milliters

Answer by Earlsdon(6098)

Remember that "milli = 1/1,000"
So, a milliliter is 1/1,000 of a liter.
and 1,000(1 milliliter) = 1,000(1/1,000 of a liter) = 1 liter.
So, 1,000 milliliters = 1 liter.
Then: 122(1 liter)= 122(1,000 milliliters) = 122,000 milliliters.
Does this make sense?